OpenStudy (loginusernamed):

If 6=p and 7=P what does 346=?

OpenStudy (anonymous):

346=p? XD

OpenStudy (anonymous):

or p=p?

OpenStudy (loginusernamed):

for 1 you find out that 67=pP so 67*x=346 then x=pP repeated till x is fufiled in Ps

OpenStudy (danjs):

probably some combo of 6 and 7, or p and P

OpenStudy (loginusernamed):

all my other questions will be for Lols

OpenStudy (anonymous):

this one makes no sense... XD

OpenStudy (danjs):

do they want you do break down the 346 into factors, and use p and P for 6 and 7 ?

OpenStudy (loginusernamed):

i gues...

OpenStudy (danjs):

like p*P = 6*7 = 42

OpenStudy (danjs):

not sure what they want, i think just ot construct 346 out of 6's and 7's, and multiples of those with any operation you choose...

OpenStudy (danjs):

and represent that with the p and P

OpenStudy (lochana):

wait

OpenStudy (danjs):

that is 462, you could just tack on a -116 onto that

OpenStudy (lochana):

yes. obviously you need to add more than p and P for the answer

OpenStudy (danjs):

no rules given, just do 346 = p+P + 333 , easy hah

OpenStudy (lochana):

found one. \[ 346 = 7*7*6 + 7*6 \] \[ 346 = 294 + 42 + 10(7-6)\] \[ 346 = 346\] P*P*p + P*p + 10(P-p)

OpenStudy (lochana):

it's ridiculous :D